Camila plans to go for vacation to Australia in 12 years from now. She estimates that she will need $21,608 for the trip. How much does she need to place in a saving account today that earns 2.24 percent per year (compounded quarterly) to accumulate this amount?
Answer #1

Let X be the amount invested today

The rate per quarter 2.24%/4 = 0.56% per quarter or 1.0056

Number of periods ( quarters in 12 years ) = 12*4 = 48 quarters  

Creating an equation

x*(1.0056)^48 = 21608

x =21608/(1.0056)48

x = 16527.3

Hence if camila places 16527.3 $ in an account earning 2.24 % per year ( quarterly compounded ) she will be able to obtain 21608 for her trip